OpenStudy (anonymous):

Solve 6x=the square root of (24 +12x). Check for extraneous solutions.

OpenStudy (ajprincess):

\(6x=\sqrt{24+12x}\) square both sides. \(36x^2=24+12x\) \(36x^2-12x-24=0\) use quadratic formula to solve for x. \[x=\frac{ -b\pm\sqrt{b^2-4ac} }{ 2a }\] compare the equation \(36x^2-12x-24=0\) with the general equation \(ax^2+bx+c=0\) and find the values of a, b nd c. then plug in them in quadratic formula and find x. Does that help? @Parden_Owens

OpenStudy (anonymous):

yes thank you @ajprincess

OpenStudy (ajprincess):

welcome:)